OpenStudy (anonymous):

PLEASE HELP WITH THIS VECTORS QUESTION! Determine the resultant: 25 N at 10° to the horizontal and 30 N horizontally. Please show all work, Thanks (:

OpenStudy (anonymous):

HEY CALLISTO REMEMBER ME? :D

OpenStudy (callisto):

Yes :) My method is a little stupid... sorry Consider the horizontal component Horizontal component of the resultant vector = 25cos10 + 30 = 54.62 Vertical component of the resultant vector = 25sin10 = 4.34 Resultant vector \(= \sqrt{54.62^2 + 4.34^2}\) = 54.79 Angle: \(tan^{-1} 4.34/54.62 = 4.53 °\) to the horizontal

OpenStudy (anonymous):

THANKS :) How should I write my therefore statement?

OpenStudy (anonymous):

Any ideas?

OpenStudy (callisto):

Eh?! Therefore, the resultant vector is 54.79N at 4.53° to the horizontal?! Hold on... seems I made mistakes there... lemme check it first :|

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Ill brb, gonna check another post real quick

OpenStudy (anonymous):

How do I check my closed posts?

OpenStudy (callisto):

Click your name -> questions asked

OpenStudy (callisto):

Alright... it should be correct :|

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

So its correct?

OpenStudy (callisto):

I think so .... I hope...

OpenStudy (anonymous):

OK THANKS ALOT! Your always a great help :) Ill be posting some more questions later on, fell free to answer :P

OpenStudy (callisto):

Okay :) Good luck!!~~~